OpenStudy (anonymous):

solve: (11/3x)-(1/3)=(-4/x^2)

OpenStudy (lgbasallote):

multiply evrything by 3x^2 11x - x^2 = -12 that should be easier now

OpenStudy (anonymous):

uhm I have no idea what that just did?

OpenStudy (alexwee123):

lgbasallote found the LCD of 3x,x,x^2 and multiplied that with everything to get rid of the denominator :o

OpenStudy (alexwee123):

then it would be a simple quadratic which you can solve easily :D

OpenStudy (lgbasallote):

lol sorry for jumping that...i assumed you knew LCMs :S

OpenStudy (mertsj):

\[\frac{11}{3x}-\frac{1}{3}=\frac{-4}{x^2}\]

OpenStudy (anonymous):

so right now the answer is 11x-x^2=-12? and it's okay but no i don't really know what i'm doing.

OpenStudy (anonymous):

thats fine with me, less work.

OpenStudy (mertsj):

Let's add the two fractions on the left and then you can cross multiply like before.

OpenStudy (mertsj):

\[\frac{11}{3x}-\frac{1}{3}\times\frac{x}{x}=\frac{-4}{x^2}\]

OpenStudy (mertsj):

\[\frac{11-x}{3x}=\frac{-4}{x^2}\]

OpenStudy (mertsj):

Cross multiply

OpenStudy (anonymous):

so it would be\[11x ^{2}-x ^{3}\div -12x\]?

OpenStudy (mertsj):

\[11x^2-x^3=-12x\]

OpenStudy (mertsj):

So let's get it equal to 0.

OpenStudy (anonymous):

oh

OpenStudy (mertsj):

\[x^3-11x^2-12x=0\]

OpenStudy (mertsj):

Now we can factor it. Factor out an x: \[x(x^2-11x-12)=0\] \[x(x-12)(x+1)=0\] x=0, x=12, x = -1

OpenStudy (mertsj):

But of course x cannot be 0 because then the denominator would be 0 so we will discard that answer and keep 12 and -1 as answers.

OpenStudy (anonymous):

Thanks. (: